A mixture of gold, zinc and mercury contains 105 g of gold, 0.662 lb of zinc and 85.0 mL of mercury (Density= 13.6 g/mL). What is the % compostion of each of the three metals in the mixture?
Answer #1

First, use mass composition, so change volume to mass

D = mass/Vol

mass = D*V = 85*13.6 = 1156 g of Mercury

mass of gold = 105 g

mass of zinc = 0.662 lb --> 0.662 *454 = 300.548 g of zinc

Total mass = 300.548+105+1156 = 1561.548 g

% Gold = 105/1561.548*100 = 6.724%

% Zinc = 300.548/1561.548*100 = 19.24%

% Merucury = 1156 /1561.548*100 = 74.02 %